OpenStudy (anonymous):

(MEDAL AND A FAN!) The volume of a cone is three hundred cubic kilometers. If the radius of the base is 5 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth.

OpenStudy (anonymous):

can someone help?

OpenStudy (anonymous):

@Nairz

OpenStudy (anonymous):

could you write it in words? you will receive two medals if you do this

OpenStudy (anonymous):

@arnavbhat r u still there?

OpenStudy (anonymous):

V=(πr^2 h)/3 so volume=pi x radius squared x height all over 3 25π x h/3=300 so 25pi x height all over 3 = 300 25πh = 900 so 25pi x height = 900 h=900/25π =11.4591559026 so the height of the cone is 11.5 km